A plane leaves an airport and flies south at 188 mph. Later, a second plane leaves the same airport and flies south at 470mph. If the second plane overtakes the first one in 2.5 hours, how long of a head start did the first plane have?
a. 2hr b. 3hr c.3.75 hr d. 4.75hr e. 6.25hr
I just can't figure this out. Please help. My e-mail address is kslau94828@aol.com
: A plane leaves an airport and flies south at 188 mph. Later, a second plane leaves the same airport and flies south at 470mph. If the second plane overtakes the first one in 2.5 hours, how long of a head start did the first plane have?
a. 2hr b. 3hr c.3.75 hr d. 4.75hr e. 6.25hr
I just can't figure this out. Please help. My e-mail address is kslau94828@aol.com

Answer by ankor@dixie-net.com(3900)

A plane leaves an airport and flies south at 188 mph. Later, a second plane leaves the same airport and flies south at 470mph. If the second plane overtakes the first one in 2.5 hours, how long of a head start did the first plane have?
:
Let t = head start time in hrs
:
When one plane overtakes the other plane. they will have traveled the same distance
Write a distance equation from this fact: Dist = speed * time
:
2nd plane travel time = 2.5 hrs
1st plane travel time = (t+2.5)
:
188(t+2.5) = 470 * 2.5
:
188t + 470 = 1175
:
188t = 1175 - 470
:
188t = 705
t = 705/188
t = 3.75 hr head start
:
:
Check the solution by confirming that each plane traveled the same distance
1st plane travel time: 2.5 + 3.75 = 6.25
:
188 * 6.25 = 1175 mi
470 * 2.5 = 1175 mi